Pls explain in detail (step by step)

Grade:12

1 Answers

Ramesh V
14 years ago

first find the resultant force....

In vector form total force on mass is : (27.5 i + 13j) N

         F = (202+152+2.20.15.cos60 )1/2 = 5(37)1/2 N

       As   F = ma

accln. = a = 6.083 m/sec2

we'll get accln, and for direction of accln,  the direction of resultant force will be same as of accln.

so direction is d = tan-1(13/27.5)

 direction of accln. = 25.3 degrees with F1

---------

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch